If the rise in the price of a bento box from $3 to $5 results in a rise in the demand for “Wakami” from 700 to 800 units, calculate the cross elasticity of demand between the two goods. State the relationship between both goods.
1
Expert's answer
2020-04-20T10:28:47-0400

The cross-price elasticity of demand is computed as:"%Change in the quantity demanded of X\/%Change in the price of Y" "%Change in the quantity demanded of X\/%Change in the price of Y"



"E_{xy}=\dfrac{\text{\% Change in the quantity demanded of X}}{\text{\% Change in the price of Y}}"



Let Wakami be good X and bento box be good Y.

If the demand for Wakami increases from 700 to 800 units as a result of increase in the price of bento box from $3 to $5, then:


"\text{\% Change in the quantity demanded of X}= \dfrac{800 - 700}{700}\times 100
\approx 14.29\%"

"\text{\% Change in the price of Y}= \dfrac{\$5- \$3}{3}\times 100
\approx 66.67\%"

Therefore:



"E_{xy}=\dfrac{14.29\%}{66.67\%}\approx 0.214"

The two goods are substitutes since their cross-price elasticity of demand is positive.